ID |
Date |
Author |
Type |
Category |
Subject |
17429
|
Fri Jan 27 11:26:31 2023 |
rana | Update | PEM | HEPA Monitor setup using WFS BLRMS | this is very exciting! Its the beginning of the task of reducing vast amounts of PEM data into human-useful (bite sized) info. Imagine if we had 60 Hz BLRMS on various channels - we would know exactly when ground loops were happening or disappearing. |
17428
|
Thu Jan 26 23:33:19 2023 |
Koji | Update | LSC | EP30-2 Epoxy bonding for Yuta | @OMC Lab
EP30-2 Epoxy bonding of V beam dumps for LSC PDs.
- Supplied 1"x0.5" glass pieces from the stock in the OMC lab.
- The black glass pieces were cleaned by IPA / Aceton / Aceton+Cotton Qtip scrub / First Contact
- 3 beam dumps are built. They require ~24 hrs to get cured.
=> Handed to Yuta on Jan 27, 2023 |
17427
|
Thu Jan 26 18:20:43 2023 |
Anchal | Update | PEM | HEPA Monitor setup using WFS BLRMS | I implemented the BLRMS on WFS1/2_I_PIT/YAW outputs and using there value, a HEPA state can be defined. I've currently set it up to use average of 10-30 Hz noise on WFS signals low passed at 0.3 Hz. For ON threshold of 100 and OFF threshold of 50, it is working for my limited testing time. To read HEPA state, one can do caget C1:IOO_HEPA_STATE. I've turned OFF HEPA for tonight's shimmer test. This is bare bones quick attempt, people can make the screen more beautiful and add more complexity if required in future. |
17426
|
Thu Jan 26 16:07:05 2023 |
yuta | Summary | PSL | PMC aligned, now PMC transmission is 0.7 | PMC aligned (Attachment #1).
Over the past month, PMC transmission is actually slowly growing from 0.68 to 0.70 (Attachment #2), since it suddenly dropped from 0.72 on Dec 27 (40m/17390). |
17425
|
Thu Jan 26 15:56:30 2023 |
Anchal | Update | ASC | 1X1 -5V sorenson tripped | [Yuta, Anchal]
Quote: |
I mounted the modified WFS boards 111B and 112B next to the whitening filter boards of existing WFS.
|
The mounting of two additional WFS demodulation boards drew too much current on -5V rail which tripped the sorenson on 1X1. This was undetected until today. Because of this, the existing WFS boards were not working either. After investicgation to beam paths and PD to board signal chain, we found out this issue. We raised the current limit on -5V supply and it came back to 5V. This brought back functioning of the exisitng WFS boards as well. We increased the current limit slightly on +5V supply too as these boards take a lot of current on +/5 V rails. But we should do this more properly by knowing what current limit the supply is set to. We'll do this part in near future after reading the manuals/wiki.
IMC WFS loops are now working. |
17424
|
Thu Jan 26 00:18:54 2023 |
Koji | Update | Cameras | Recording CCD cameras | Connect any video signal supported by the adapter. Use Windows / Mac or any other OS. If it keeps complaining, contact Magwell support. |
17423
|
Wed Jan 25 18:01:21 2023 |
yuta | Summary | BHD | BH44 setup isolated, but 60 Hz noise stays the same | [Paco, Anchal, Yuta]
Isolating BH44 setup from the rest didn't help reducing the 60 Hz noise.
Frequency noise from IMC also seems unchanged before and after BH44 installation.
Isolating BH44:
- To see if BH44 setup installed is causing the 60 Hz issue, we compared the spectra of FPMI sensors with BH44 setup and with BH44 setup disconnected.
- In the latter configuration, BH44 setup was isolated from the rest by disconnecting the SMA cables and the RFPD power cable, as shown in Attachment #1.
- There was no significant difference in the spectra with BH44 and with BH44 isolated.
We have even put the old AS156 IQ demodulator board we have pulled out to insert BH44 IQ demodulator board back, but didn't change.
- We have also disconnected the 22 MHz generation setup around 40m Frequency Generation Unit at 1X2 for switchable IMC/AS WFS, but it also didn't help.
- Attachment #2 is the orignal spectra with both arms locked with POX and POY, feeding back to respective ETMs (MICH is not locked), and Attachment #3 is those with BH44 setup isolated, AS156 IQ demod back, and 1X2 22MHz generation isolated. Both look basically the same.
- BH44 setup was reverted after the comparison.
IMC frequency noise:
- As adding a resonant gain at 60 Hz helped reducing the 60 Hz noise (40m/17419), the noise might be from frequency noise. It also explains why it is not present in MICH when ETMs are mis-aligned, and only present when one of the arms is involved (40m/17413).
- To see if the frequeny noise at 60 Hz increased after BH44 installation, I compared the spectrum of C1:IOO-MC_F_DQ on January 11 (same Wednesday) with that measured today at almost the same time.
- Attachment #4 is the result. 60 Hz noise and its harmonics seems almost the same in MC_F. It is rather noisy today in other frequencies, but not at 60 Hz.
Next:
- Read the book. |
17422
|
Wed Jan 25 16:58:19 2023 |
Alex | Update | Cameras | Recording CCD cameras | Thus far, the software needed for the Magewell video encoder has been successfully installed on Donatella. OBS studio has also been installed and works correctly. OBS will be the video recording software that can be interfaced via command line once the SDI video encoder starts working. (https://github.com/muesli/obs-cli)
So far, the camera can not be connected to the Magewell encoder. The encoder continues to have a pulsing error light that indicates "no signal" or "signal not locked". I have begun testing on a secondary camera, directly connected to the Magewell encoder with similar errors. This may be able to be resolved once more information about the camera and its specifications/resolution is uncovered. At this time I have not found any details on the LCL-902K by Watec that was given to me by Koji. I will begin looking into the model used in the 40 meter next. |
17421
|
Wed Jan 25 14:33:07 2023 |
rana | Summary | BHD | REFL55 visually inspected, BH44 Kapton taped | Kapton tape is a good insulator, so its a good block for 60 Hz. But it is mostly useless for RF since the capacitance between the mount and the table is
C = epsilon * A / d
For Kapton the dielectric constant is 3.5, the PD mount area is 10 cm x 10 cm, and the film thickness is ~50 um. So C ~ 5 nF.
Z ~ 1 / (2 * pi * 44 MHz) / C
~ 0.5 Ohms |
17420
|
Wed Jan 25 12:49:14 2023 |
Radhika | Update | ALS | XARM green laser lock debugging |
I returned the half-wave plates on the XEND table back to their original angles, and restored the loop configuration with the PDH servo box. I returned the PD gain to 40 dB (original setting), and set the servo gain knob to 6. This was the region of highest loop stability, with the lock holding for a few seconds (as before). The control signal on the scope did not look intuitive - the peaks of the control signal corresponded with zero crossings of the error signal.
Paco encouraged me to retake transfer function measurements of the PDH servo box. The main takeaway is the PDH servo (boost on) has the expected frequency response at a gain setting of 3 or under, up to 100 mVpp of input. Attachment 1 shows the frequency response at a servo gain of 2, for varying input amplitudes.
The rest of the bode plots correspond to servo gain of 4, 6, 8, and 10 (boost on). The saturation LED would turn on above a gain value of ~3.25, so these results can't be analyzed or interpreted. But it does seem like a steep, low-frequency jump is a signature of the saturated servo. This jump doesn't appear with 10 mVpp input, at least at or above 1 Hz. |
17419
|
Wed Jan 25 11:35:20 2023 |
yuta | Summary | BHD | REFL55 visually inspected, BH44 Kapton taped | Following was done to investigate 60 Hz noise issue, but no significant change in the FPMI noise observed.
REFL55 inspection:
- Even before BH44 installation, we have been experiencing flaky REFL55 RF output. When some work was done at AP table or something, sometimes the amplitude of REFL55_I and REFL55_Q goes very low, and/or offset changes. This was usually fixed by touching the RF output of REFL55.
- So, we took out REFL55 and opened the back lid to inspect. RF output seemed rigid and the SMA connector was properly grounded to the box; didn't find any issue (Attachment #1).
- REFL55 was put back to its original position, and the cables were also put back.
BH44 Kapton tape:
- I realized that other RFPDs have Kapton tape in between the RFPD golden box and the black mount, but not for BH44 we recently installed.
- I have checked that the golden box of BH44 and the optical table is not grounded when RF output and the DB15 cable was disconnected, but is gounded when they are connected, just like BH55.
- Anyway I removed BH44 and put a Kapton tape (Attachment #2), just in case, and BH44 was put back to its original position, and the cables were also put back.
FPMI noise spectra after the work:
- Attachment #3,4,5 are noise spectra of FPMI BHD sensors when FPMI is RF locked with AS55_Q, REFL55_I, and REFL55_Q, and LO_PHASE is locked with BH55 with the following configurations.
- Attachment #3: whitening/unwhitening filters for AS55, REFL55, POX11, POY11, BH55, BH44 turned on (nominal configuration after lock acquisition)
- Attachment #4: whitening/unwhitening filters for AS55, REFL55, POX11, POY11, BH55, BH44 turned off. No significant change except for expected whitening filter transfer function.
- Attachment #5: whitening/unwhitening filters for AS55, REFL55, POX11, POY11, BH55, BH44 turned on, 30 dB resonant gain at 60 Hz, Q=10 in CARM loop. Significant 60 Hz reduction everywhere. This was not observed when resonant gain at 60 Hz was put in DARM loop (only 60 Hz at AS55_Q was reduced). 60 Hz noise mainly coming from something in CARM loop?
Don't forget to:
- Put a beam dump for BH44 |
17418
|
Wed Jan 25 10:02:47 2023 |
yehonathan | Update | | Earthquake, MC3 watchdog tripped | We came this morning and noticed the FSS_FAST channel was moving very rapidly. Short inspection showed that MC3 watchdog got tripped. After reenabling the watchdog the issue was fixed and the MC is stable again.
There is a spike in the seismometers 8 hours ago and it was probably due to the 4.2 magnitude earthquake that happened in Malibu beach around that time. |
17417
|
Tue Jan 24 21:29:31 2023 |
yuta | Summary | BHD | Electronics diagram around BH44 and BH55 | 1) Turning the whitening filter before the ADC on/off didn't changed the relative height of 60 Hz peak compared with the noise floor. When the whitening filter was turned on, increase of the dark noise measured at C1:LSC-****_(I|Q)_IN1 was roughly consistent by eye with the whitening filter transfer function (gain of 1 at DC, ~15 Hz zero x2, ~150 Hz pole x2), which suggests the 60 Hz pickup is before the whitening filter.
2) Attached is the electronics diagram around BH44 and BH55. |
17416
|
Tue Jan 24 21:04:59 2023 |
Anchal | Update | ASC | AS WFS path beam profiled | I completed the mode matching calculation today and found good solution with 360.6 mm ROC PLCX lens at -1.2 m from z=0 point. I placed the lens there today and aligned all mirrors to get centered beam on both WFS PDs when the flipper mirrors are flipper up. This alignment would probably require tweaking everying we flip the mirrors as the flipper mirrors do not come back to same position usually.
I mounted the modified WFS boards 111B and 112B next to the whitening filter boards of existing WFS. Now to switch over, onewould need to transfer the 8 RF lemo cables and the 2 IDE ribbon cables.
I'm working on rtcds model to read AS WFS data and handle it separately. I'll keep a WPICS binaruy switch to switch between IMC WFS or AS WFS. I need to figure out some build issues on this work still.
|
17415
|
Tue Jan 24 16:12:07 2023 |
rana | Summary | BHD | FPMI displacement noise with 60 Hz harmonics side lobes | 1) do a comparison with the whtiening before the ADC on/off. This will tell us if it is pickup before the whtiening filter or not.
2) If there are ground loops made by 44 MHz setup, we want to draw a simple diagram which includes which sides are grounded and which have transformers. How about make a drawing to bring to the group meeting tomorrow? IN the lab we have these coaxial BALUNs for making a 1:1 transformer coupling.
3) Another source of 60 Hz is the unintentional demodulation of spikes made by the Sorensen switching supplies: they produce spikes all the way up to 100 MHz, so if they land near 44 MHz, you may get some 60 Hz on the demodulation. You should be able to see this with a dipole antenna or a hoop antenna. |
17414
|
Tue Jan 24 11:21:59 2023 |
yuta | Summary | BHD | FPMI displacement noise with 60 Hz harmonics side lobes | Just to show how bad 60 Hz noise is, I compared FPMI displacement noise with pre-BH44 era (measured on Jan 13, 40m/17400).
Blue curve in Attachment #1 is the sensitivity with FPMI locked with RF in pre-BH44 era, and pink curves are that measured today (C1:CAL channels are currently unavailable due to 0x2000 appeared after running restatAllModels.sh).
60 Hz + harmonics pedestals are apparent today, but was not there on Jan 13. Today, DARM could be handed over to AS55_Q from POX11-POY11, but CARM could not be handed over to REFL55_I from POX11+POY11 (this was possible last night).
Attachment #2 shows FPMI error signals when electronic FPMI is locked. Too much 60 Hz, especially in REFL55_I_ERR and AS55_I_ERR (note that REFL55_Q is used for MICH lock, but AS55_Q is not in-loop yet when this screenshot was taken.)
Next:
- Fix c1cal 0x2000 issue
- Fix REFL55 loose RF output
- Disconnect cables in BH44 to open possible ground loops made during BH44 installation (especially 44 MHz generation part??). |
17413
|
Mon Jan 23 22:51:17 2023 |
yuta | Summary | BHD | 60 Hz harmonics side lobe investigations | [Paco, Yehonathan, Yuta]
Since we have installed BH44, we are seeing side lobes of 60 Hz + harmonics in AS55, REFL55, BH55, BH44, preventing us from locking FPMI BHD (40m/17405).
BH55 RF amp removed:
- We have noticed that the side lobes are there in BH55 (but not in BH44) when LO-ITMX single bounce is fringing (ETMs and ITMY mis-aligned).
- Changing whitening gains and turning on/off whitening/unwhitening filters didn't help.
- When LO-ITMX single bounce is locked with BH55, the side lobe in BH55 reduces.
- Dithering LO1 at 11 Hz created 180 +/- 11 Hz signal, which confirms that this side lobes are from the up conversion of optic motion.
- We thought it could be from RF saturation, so we have put a 55-67 MHz bandbass filter (SBP-60+) in between BH55 RFPD and RF amp (ZFL-1000LN+; 40m/17195). Didn't help.
- We then removed the RF amp. This largely reduced the side lobes (but still some at 180 Hz). We could lock LO-ITMX single bounce without the RF amp, so we decided to remove it for now.
Side lobes only when one of the arms is locked:
- When ETMs are mis-aligned, MICH fringing and BHD fringing, there are 60 Hz + harmonics, but the side lobes are not there.
- But with Xarm is locked (ETMY, ITMY mis-aligned) or Yarm is locked (ETMX, ITMX mis-aligned), the side lobes appear in AS55, REFL55, BH55, BH44.
- Changing whitening gains and turning on/off whitening/unwhitening filters didn't help.
- As the error signals are normalized by TRX and TRY, we turned on/off the power normalization, but didn't help.
- Switching 60 Hz comb in BS, ITMX, ITMY, ETMX, ETMY suspension damping didn't help.
POY11 Investigations:
- When ETMs are mis-alined, POX11 had relatively large 60 Hz + harmonics, but almost none in POY11 (unlike other RFPDs; see Attchment #1).
- However, when ETMY is aligned and Yarm is loked with POY11, the side lobe grows in POY11.
- Changing the feedback point from ETMY to ITMY or MC2 didn't help.
- We have unplugged the IQ demod board for BH44 from the eurorack (without removing the cables) and removed the fuse for the power supply of the RF amp for 44 MHz generation (40m/17401), but these also didn't help.
- We have also tried locking Yarm with REFL55(= ~2 x POY11), BH55(= ~10 x POY11), ALSY(= ~2000 x POY11) , but the side lobes were always there.
Next:
- Disconnect cables in BH44 to open possible ground loops made during BH44 installation (especially 44 MHz generation part??).
- Check if the noise was there before BH44 installation using past data. |
17412
|
Mon Jan 23 20:50:58 2023 |
Anchal | Update | ASC | AS WFS path beam profiled | I measured the expected beam profile by WFS photodiodes by measuring the beam when mode cleaner was unlocked from the point where beam is picked for WFS. See attachment 1 for beam details. z=0 is the point in the path where AS beam will merge.
For measuring the beam profile of AS beam, I had to focus it using a lens. I picked up a 360.6 mm ROC lens and placed it at z=-67 inch point. Then I profiled the beam at some comfortable section of the path and fitted it. with reverse z-axis. Using this method, I can place the lens back and obtain the original beam back. Attachment 2 shows this fitting process and identification of the original beam profiles. I plotted the AS beam profiles again in attachment 3 and saved them for seeding mode matching effort later. Note that we don't want to be super accurate here, so I did not do any error analysis, just wanted to finish this fast. Also pardon me for the bad quality plots, I did not want to learn Matlab plotting to make it beautiful.
Note: There is significant astigmatism in both IMC reflection beam and AS beam. This could be due to beam going through far off-center on lens. Something to keep in mind, again this measurement is not ideal in terms of precision but this large an astigmatism could not be due to measurement error.
Next:
- Identify correct len(s) and their positions
- Align the AS beam to WFS heads
- Test the full signal chain.
|
17411
|
Mon Jan 23 16:31:23 2023 |
rana | Update | ALS | DFD and Phase tracker AM coupling | Both the TF measurement and the noise measurements are useful, but the nosie measurement is much more meaningful. Since we expect the main coupling to be incoherent, what we really want is a noise budget style measurement:
- Measure the FM noise spectrum with only a single sine wave into the Moku.
- Same as #1, but with the AM noise added as you already did.
- Estimate the noise budget contribution by doing PSD subtraction, and then scale that by the excitation magntiude. This will be the contribution of beat amplitude noise to the measured calibration.
Quote: |
I took transfer function measurement of DFD AM coupling using noise excitation.
|
|
17410
|
Mon Jan 23 11:20:44 2023 |
JC | Summary | BHD | BH44 RFPD optical path and LO/AS camera | Here's the beam path of BH44. |
17409
|
Sat Jan 21 18:01:06 2023 |
Anchal | Update | ALS | DFD and Phase tracker AM coupling | I took transfer function measurement of DFD AM coupling using noise excitation.
Noise excitation setup
Noise is injected using C1:BAC-SPARE_CH14_EXC using awggui which is filtered by a foton filter to simulate the real beatnote RF amplitude noise measured by taking quadrature sum of C1:ALS-BEATY_FINE_I_OUT and C1:ALS-BEATY_FINE_Q_OUT. See attachment 1.
The DAC output is connected to MP1 at CH1. MP1 is set to run in waveform generation mode with following settings:
- Carrier frequency: 45 MHz
- Carrier Level: 500 mVpp
- No offset or phase offset
- Amplitude modulation ON
- Modulation slope: 100%/V
- Source Input: Ch1
The AWGUI is set to excite C1:BAC-SPARE_CH14_EXC using settings mentioned in attachment 2.
With this setup, the RF amplitude noise is simulated with MP1 and DAC excitation.
Transfer function measurement
With AWGGUI running as mentioned above, I simply used diaggui in spectrum mode for channels C1:BAC-SPARE_CH14_EXC and C1:ALS-BEATY_FINE_PHASE_OUT_HZ. The second channel is already calibrated into Hz, and the first channel is in counts. To convert it into voltage of amplitude fluctuation, I first converted DAC excitation to voltage by assuming 16 bit DAC with +/- 10 V range, this gives conversion constant of 10/2**15 V/cts. Then since MP1 is doing 100%/V AM modulation, for 500 mVpp RF level, this means 0.25 V/V AM modulation. Multiplying these two together gives, 7.6294e-5 V/cts. I put this number in teh diaggui calibration for C1:ALS-BEATY_FINE_PHASE_OUT_HZ.
This created the transfer function measurement attached in attachment 3.
The measurement resulted in roughly 2kHz/V AM to frequency coupling in DFD + phase tracker setup. The previous measurement with coherent sinusoidal excitation was exactly a factor of 1000 less than this, so I believe I might have made some error in calibrating or there could be an error in the previous elog. Please check my calculations. But a solid thing to note is the coherence measured below 1Hz. I'll do more sophisticated analysis on weekdays.
I also think that coherence was low because of low excitation. We should redo this test with more noise power to get good coherence in all frequency band to have good idea of what would happen to ebatnote RF amplitude noise at all frequencies.
Mon Jan 23 11:47:23 2023 Adding Attachment 4:
I realised that with the noise excitation setup set to mimic real beatnote amplitude noise with very low frequency noise as it is seeded with Moku:Pro, the measured frequency noise by the DFD+Phase Tracker setup at C1:ALS-BEATY_FINE_PHASE_OUT_HZ is an indicator of how much RF amplitude noise of beatnote contribute to the frequency noise measured by DFD+Phase tracker. Attachment 4 is the spectrum measured during this measurement. |
17408
|
Sat Jan 21 15:32:40 2023 |
Anchal | Update | ASC | AS WFS path nominally set | I've completed the beam redirection path for AS beam to WFS heads in a nominal way. By that I mean that all mirrors (M1, M2, M3, and M4) are now in their final positions and we will need to install one or two lenses to collimate the beam to match the mode that the WFS path is expecting as it has it's on the focusing lens before the photodiodes. For this last part, I think the fasted way would be to profile the beam and calculate the correct lens and position rather than trial and error as the beam intensity is very low for estimating the beam size by eye.
IMC WFS state: Flip M1 and M2 down.
AS WFS state: Flip M1 and M2 up. |
17407
|
Fri Jan 20 20:13:20 2023 |
Anchal | Update | ASC | Installed 2 flipper mirrors for handingl MC reflection beam to camera | After discussions with Yuta, I figured that a better optical layout is possible which does not interfere with the existing IMC WFS path at all. So I reset the IMC WFS path today (and zeroed RF offsets again) and changed the MC reflection camera and MC reflection beam dump (black hole) position to create space for a flipper mirror that will pop up in the IMC WFS path and steer in the AS beam. New proposed path is shown in the photo in cyan. Red is MC reflection beam, yellow is IFO reflection beam and orange is teh AS beam that we will pick up using flipper mirror M1. Note that I found an intense 6.4 mW ghost beam coming out of the interferometer in between IFO refl and MC refl beams. This beam is shown in pink which I have dumped now. This beam was earlier not dumped. We will need to investigate more on the source of this beam and correct it in the next vent. |
17406
|
Thu Jan 19 20:35:54 2023 |
Anchal | Update | ASC | Installed 2 flipper mirrors for handingl MC reflection beam to camera | Today I installed two flipper mirrors M3 and M4 (see attached photo) to create alternate route for MC reflection camera beam. Both these mirrors are Y1-1037-45S. In nominal operation where IMC is using the WFS, we will keep M3 upright and M4 flipped down. When using WFS for AS, M3 will be flipper down and M4 will be upright to save the camera from the high intensity MC reflection beam.
Note that everytime M3 is flipper and put back upright, the alignment into WFS would need to be tuned as the flipper apparatus does not come back to same alignment everytime. I centered the beams on the WFS heads today and zeroed RF offsets usingC1IOO_WFS_MASTER>!Actions>Correct WFS RF Offsets script. After this, the IMC WFS loops are working as expected atleast for last 15 minutes that I have monitored them. Hopefully, this will remain consistent.
Upcoming work:
- Change the steering mirror that steers the beam to black hole to be a flipper mirror too as AS beam strength (measured when MICH was locked to bright port) is 0.3 mW and IMC WFS heads combined power is 0.5 mW in nominal operation, so we can not afford to dump any AS beam light.
- Put flipper mirror M1 and fixed mirror M2 mentioned in 40m/17320 for steering AS beam to IMC WFS heads.
|
17405
|
Thu Jan 19 18:15:48 2023 |
yuta | Summary | BHD | BH44 RFPD optical path and LO/AS camera | [JC, Paco, Anchal, Yuta]
- We installed the new BH44 beam path and RFPD.
- JC installed the new beam path for the LO/AS camera.
- We succeeded in locking LO Phase with BH44_Q_ERR, but didn't attempt FPMI BH44 because we noted large 60 Hz harmonics in most of our RF error signals.
BH44 RPFD/Camera installation:
- We picked off LO/AS beam path previously going to the camera, and installed a Y1 (45 deg, s-pol) mirror, a f=150mm lens and the RFPD (Attachment #1). We initially tested it using the incandescent light from a flashlight and then aligned the beam, we also made sure it's not saturating.
- Using the spurious transmission from the mirror mentioned above, we steered a new beam path for the camera and realigned it using another short focal length lens (f ~ 100 mm).
LO Phase control:
- We increased the whitening gain from 0 dB to 42 dB for both C1:LSC-BH44_I and C1:LSC-BH44_Q, and zeroed the offsets. Even before this step we could see a fringe from BH44, which is quite promising!
- After alignment was recovered on the LO/AS path, we succeeded in locking the single bounce (ITMX) LO phase using BH44_Q. Here the configuration was FM4, FM5 and a gain of ~ 5 * 0.5 = 2.5 (to match the typical BH55_Q error point).
- While BH44_Q was used to control the LO phase, we saw the BH55_Q was not zero but actually almost at max fringe value (see Attachment #2). This implies the BH44_Q is indeed orthogonal to BH55_Q with respect to the LO Phase!
FPMI lock:
- We locked electronic FPMI but noted a large 60 Hz + harmonics component in the RF error signals including AS55, BH55, REFL55, and BH44 (see Attachment #3). We could hand off to FPMI and even locked the LO phase with BH44_Q, but we were not sure the BHD_DIFF error signal was fit for handoff to FPMI BHD. Therefore we stopped here.
60 Hz + harmonics:
- We did a quick investigation around the areas we have been working in the lab to see if we had introduced this noise in any obvious way. First we checked the new amplifier for the 44 MHz LO, we briefly removed its power but the 60 Hz noise remained. Then we checked the AP table, but nothing had really changed there. We also disconnected and removed the rolling cart with the marconi and other instruments from the LSC rack. Finally, we turned all the lights down. None of these quick fixes changed the amount of noise.
- We also tried looking at these error signals under different IFO alignment and feedback configurations. We always see the noise in the AS55 and REFL55 quadratures, but not in BH44, BH55 or BHD_DIFF unless MICH is locked.
Next steps:
- Investigate more into 60 Hz noise, why? where?
- Measure sensing matrix with LO Phase locked with BH44 and BH55 to make comparison.
- FPMI-BH44 |
17404
|
Thu Jan 19 14:58:40 2023 |
rana | Summary | BHD | IQ demod orthogonal | the problems with these circle plots:
- you have to make the aspect ratio of the PDF correct or else it doesn't look like a circle
- what we care about is the deviation from circle, so you should plot the difference in a way that let's us see the difference, not just that it sort of looks like a circle. This is similar to how we plot the residual when doing fits.
|
17403
|
Thu Jan 19 12:12:09 2023 |
Anchal | Summary | BHD | IQ demod orthogonal | By sending two frequencies offset from eachother to LO input and RF input, we measured the remaining phase difference between I and Q outputs of this demod board and correct that by setting C1:LSC-BH44_PHASE_D to 86.2 degrees and balancing the amplitudes by putting C1:LSC-BH44_Q_GAIN to 1.747. See attached for XYPlot after correction. |
17402
|
Wed Jan 18 20:57:53 2023 |
Paco | Summary | BHD | IQ demod orthogonal | I tested a spare IQ demod board labeled 33.3 MHz (closer to 44 MHz than the 165 MHz we had started using) and using the Moku adjusted the trim caps after the transformer T1 to adjust orthogonality (using an oscilloscope). The orthogonality seems quite good on this board and it seems to be working fine, so I decided to swap out the BH44 (previously AS165) IQ demod board for this one. To do this I first unpowered the amplifier, then disconnected the load (IQ demod board) then release from the Eurocrate, then add the new board.
Finally, using Marconi at 11.066195 * 4 to get close to the 44 MHz LO frequency, I measured a 63.9 Hz tone at the C1:LSC-BH44_I_IN1 and C1:LSC-BH44_Q_IN1 channels (whitening gain 0 dB). The measured angle between these two channels was 86.97 deg, so the orthogonality is much better now. The gain imbalance is also relatively better, Q/I ~ 0.57 |
17401
|
Tue Jan 17 20:03:19 2023 |
yuta | Summary | BHD | BH44 installations: IQ demodulator is not orthogonal | [Anchal, Paco, JC, Yuta]
We have started hardware installations for BH44 RF PD. 44 MHz LO generation and signal chain from IQ demodulator was checked successfully, but found that IQ demodulator is not orthogonal.
RF PD Interface:
- We have unplugged Ch2 of the RFPD Interface (labeled "Special") to re-use it for BH44. Ch2 was used for "UNIDENTIFIED" RFPD. DB15 cable was routed to ITMY table and connected to BH44 RF PD (40m/17398) now sitting on the cover of ITMY table. See Attachment #1.
- Finding a DB15 RF PD interface cable was easy because of the organization work!
44 MHz LO generation:
- LO for BH44 was generate following the scheme proposed in 40m/17319.
- 11 MHz LO from RF distributor labeled "+16 dBm" (measured to be 16.5 dBm) and 55 MHz LO labeled "SPARE 55" (measured to be 2.26 dBm) was mixed with a mixer ZFM-1H-S+ (using 11 MHz as LO, and 55 MHz highpass filtered with SHP-50+ as RF). The mixer output was lowpass filtered with SLP-50+, and amplified with ZFL-500LN+, which gave 8.07 dBm at 44 MHz. The second heighest peak was -11.53 dBm at 22 MHz, which seems low enough. See Attachment #2 for the photo of the setup.
IQ demodulation:
- We have used unused IQ demodulator labeled "AS165" to use it for BH44. See Attachment #3.
- We have quickly checked if the IQ demodulator is working or not with LO from BH55, PD input of 55 MHz generated using Moku to see I and Q outputs. The outputs are sine waves at frequency consistent with the difference between LO frequency and "PD input" frequency, and the phase was off as expected. Q output was ~4 dBm higher than I output.
Measured diff of BH44:
- After CDS modifications where done, BH44 IQ demodulator was tested by using 44 MHz LO generated in a method mentioned above, and injecting 11.066195 * 4 MHz signal from Moku as PD input. This gave ~75 Hz signal in C1:LSC-BH44_I and C1:LSC-BH44_Q.
- With 0dB whitening gain and whitening/unwhitening filters off, gain imbalance was measured to be Q/I=137.04/62.49=2.19, and measured phase difference to be PHASE_D=27.21 deg (see Attachment #4; gpstime=1358051213).
- With 0dB whitening gain and whitening/unwhitening filters on, gain imbalance was measured to be Q/I=138.44/63.21=2.19, and measured phase difference to be PHASE_D=26.95 deg (see Attachment #5; gpstime=1358051325138). This is consistent with whitening/unwhitening off, and noise is smaller, which mean whitening/unwhitening filters are probably working fine.
- IQ demodulator board might be not working properly, as I and Q signals are not quite orthogonal.
Model changes:
- We modified c1lsc, c1hpc and c1cal model for BH44.
- BH44 ADC pins were identified and connected for RFPD phase rotator.
- The signals are sent to c1hpc through IPC where BH44 is now available for feedback loops in single and dual demodulation.
- The whitening filter controls and anti-aliasing filter enable buttons were created in c1iscaux slow machine db files.
- MEDM screens are updated accordingly (see Attachment #6).
Next:
- Use different IQ demodulator board that has better IQ orthogonality.
- Connect BH44 RF PD and use 44 MHz test input to check the signal chain.
- Install BH44 RF PD optical path.
- Try locking LO_PHASE with BH44. |
17400
|
Fri Jan 13 18:54:59 2023 |
yuta | Summary | LSC | Investigations of LO phase locking in FPMI BHD | [Paco, Yuta]
After several hours of unattended IFO, we realigned the IFO and somehow BH55_Q error signal looked better, LO_PHASE locking was more robust, and we could lock FPMI BHD.
Sensing matrix:
- Attached #1 is the sensing matrix with FPMI BHD locked, with LO_PHASE locked with BH55_Q using AS4, and below is the calibrated one.
Sensing matrix with the following demodulation phases (counts/m)
{'AS55': -168.5, 'REFL55': 92.32, 'BH55': -110.0}
Sensors DARM @307.88 Hz CARM @309.21 Hz MICH @311.1 Hz LO1 @315.17 Hz
AS55_I (+0.43+/-2.22)e+11 [90] (+5.26+/-1.26)e+11 [0] (-0.15+/-5.88)e+10 [0] (+0.05+/-1.94)e+09 [0]
AS55_Q (-3.73+/-0.19)e+11 [90] (-0.09+/-1.06)e+11 [0] (+0.21+/-6.82)e+09 [0] (-0.24+/-2.41)e+08 [0]
REFL55_I (+0.08+/-1.17)e+12 [90] (+3.14+/-0.07)e+12 [0] (-0.01+/-1.31)e+11 [0] (+0.07+/-1.03)e+09 [0]
REFL55_Q (+0.51+/-1.09)e+10 [90] (+1.80+/-2.35)e+10 [0] (+0.01+/-1.44)e+09 [0] (-0.79+/-5.41)e+07 [0]
BH55_I (+1.20+/-0.34)e+11 [90] (+0.26+/-1.07)e+11 [0] (-0.04+/-1.33)e+10 [0] (-1.07+/-1.82)e+09 [0]
BH55_Q (+1.12+/-1.52)e+11 [90] (-3.40+/-1.98)e+11 [0] (-0.01+/-4.16)e+10 [0] (-5.55+/-1.66)e+09 [0]
BHDC_DIFF (+6.95+/-0.25)e+11 [90] (+0.01+/-1.60)e+11 [0] (-0.70+/-4.49)e+09 [0] (+4.24+/-4.14)e+08 [0]
BHDC_SUM (+7.06+/-2.64)e+10 [90] (-0.32+/-3.24)e+10 [0] (+0.05+/-4.26)e+09 [0] (+1.30+/-6.33)e+08 [0]
BH55_Q:
- We are not sure why BH55_Q error signal got better. Peak to peak amptidue of BH55_Q when LO_PHASE is not locked is at around 800 (even if LO_PHASE locking cannot be stably locked), but when we couldn't lock LO_PHASE stably, it was noisier. This suggests that BHD alignment is not bad.
- Attachment #2 shows the spectrum of BHDC_DIFF, BH55_Q (measured at LO_PHASE_IN1) and AS55_Q when FPMI is locked with AS55_Q, and LO_PHASE is locked with BH55_Q. Dashed darker curves are when we could not lock LO_PHASE stably. You can see broad noise in BH55_Q at around 60 Hz and its harmonics when it is noisy, but they are not present when LO_PHASE can be locked stably. Also, AS55_Q stays the same, which suggests that the cause is not in FPMI to AS55, but in BHD path to BH55.
Sensitivity comparison with different LO_PHASE locking actuator:
- We compared FPMI sensitivity with LO_PHASE locked with LO1 (red) vs with AS4 (orange). Didn't change, as expected (Attachment #3).
Next:
- Investigate what is causing noisy BH55. Is it FPMI alignment, BHD alignment, suspensions in LO/AS paths, or electronics?
- Try locking LO_PHASE with an offset to BH55_Q to see if BHD_DIFF sensitivity to DARM changes, and to see if we can reduce BHD_DIFF dark noise contribution to FPMI sensitivity
- Try BH55+audio and BH44 to lock LO_PHASE.
- PRMI and PRFPMI |
17399
|
Fri Jan 13 14:20:34 2023 |
yuta | Summary | LSC | Calibration friendly FPMI BHD | [Paco, Yuta]
Gains in DARM are corrected to make it more calibration friendly.
DARM error signals
- 0.19 * POX11_I - 0.19 * POY11_I
- 1 * AS55_Q
- -0.455 * BHD_DIFF (needs to be checked if LO phase is different)
DARM gain:
- C1:LSC-DARM_GAIN = 0.04 (it was 0.015 to have UGF of ~150 Hz)
Online calibration:
- FM2 for C1:CAL-DARM_CINV was turned on, which is a calibration for AS55_Q in FPMI; 1 / (3.64e11 counts/m) = 2.747e-12 m/counts (see sensing matrix below; consistent with 40m/17369).
- FM2 for C1:CAL-DARM_A was updated to 10.91e-9 (40m/16977).
- C1:CAL-DARM_W_OUT will be our calibrated FPMI displacement in meters. This is correct with BHD_DIFF locking, if the BHD_DIFF is balanced with AS55_Q before DARM_IN1.
FPMI sensitivity:
- Attached plot shows the sensitivity of FPMI with AS55_Q and BHD_DIFF, plotted together with their dark noise.
- The sensisitivty was measured with calibration lines off and notches off, which removed the forest of lines we saw on Jan 11 (40m/17392).
FPMI sensing matrix:
- Attached is a screenshot of uncalibrated sensing matrix MEDM screen. Audio demodulation phase for DARM was tuned to have stable sign.
- The following is calibrated sensing matrix measured today with FPMI locked with AS55_Q. BH55 and BHD_DIFF have large uncertainties because LO_PHASE locking is not stable.
Sensing matrix with the following demodulation phases (counts/m)
{'AS55': -168.5, 'REFL55': 92.32, 'BH55': -110.0}
Sensors DARM @307.88 Hz CARM @309.21 Hz MICH @311.1 Hz LO1 @315.17 Hz
AS55_I (-3.28+/-0.90)e+11 [90] (-0.05+/-2.53)e+11 [0] (+0.47+/-2.14)e+10 [0] (-0.06+/-1.76)e+09 [0]
AS55_Q (-3.64+/-0.08)e+11 [90] (-0.09+/-8.30)e+10 [0] (-0.28+/-2.24)e+09 [0] (+0.12+/-1.16)e+08 [0]
REFL55_I (+0.07+/-1.24)e+12 [90] (+3.32+/-0.06)e+12 [0] (-0.01+/-1.39)e+11 [0] (+0.00+/-1.12)e+09 [0]
REFL55_Q (-0.98+/-2.46)e+09 [90] (-4.68+/-2.04)e+09 [0] (+0.08+/-1.00)e+09 [0] (+1.73+/-5.69)e+07 [0]
BH55_I (-6.84+/-2.47)e+11 [90] (+0.43+/-2.32)e+11 [0] (-0.33+/-6.21)e+10 [0] (-2.51+/-8.09)e+09 [0]
BH55_Q (+5.68+/-1.57)e+11 [90] (-0.17+/-2.52)e+11 [0] (-0.46+/-4.39)e+10 [0] (+0.73+/-5.01)e+09 [0]
BHDC_DIFF (-2.48+/-3.47)e+11 [90] (-0.09+/-1.99)e+11 [0] (+1.56+/-4.11)e+10 [0] (-0.49+/-6.78)e+09 [0]
BHDC_SUM (-2.12+/-1.84)e+10 [90] (+0.35+/-1.44)e+10 [0] (-1.30+/-4.18)e+09 [0] (-0.03+/-8.17)e+08 [0]
Current status:
- After locking FPMI BHD to get the FPMI sensitivity today, we are struggling to re-lock again. LO_PHASE locking is glitchy today for some reason. To be investigated.
|
17398
|
Fri Jan 13 13:34:12 2023 |
Anchal | Summary | BHD | BH44 tuned and transimpedance measured | I've tuned one gold box RFPD to be resonant at 44.26 MHz and I left the notch to be near 66 MHz, however, it is only effective by 10 dB. Attached is the measured transimpedance using the test port. This measurement should be updated with PD testbed measurement.
This photodiode is ready to be installed for the dual RF lO phase locking scheme.
Thu Jan 19 15:06:43 2023 Updating the measurement with Moku:Pro calibration TF |
17397
|
Thu Jan 12 18:51:17 2023 |
Paco | Update | ALS | DFD and Phase tracker AM coupling | [Anchal, Paco]
We measured DFD AM coupling; it seems to be minimum at higher RF input and low modulation depths as expected.
To do this, we set up Moku:Lab for AM ext with a spare DAC channel (C1:BAC-SPARE_CH14_OUT) which we send a swept sine excitation using diaggui. We vary the carrier level, and the modulation depth (every time we changed the level, we run the phaseUGF.py script to allow the phase tracker to adjust its loop gain properly). Attachment #1 shows the results, showing the finite bandwidth effect of the phase tracker as well as the mean magnitudes of the AM coupling below 100 Hz. This measurement and the script live in
/opt/rtcds/caltech/c1/Git/40m/measurements/ALS/DFD_calibration/
What this means for ALS calibration:
It seems the residual AM coupling for a typical RF input level from our ALS beatnote corresponds to the couplings of ~ 2 Hz/V. This means that if the RF input level is fluctuating by 100mV, our residual beat frequency fluctuations only move by 200 mHz. This is not the case when the arms are locked... there the beat level stability is closer to 1 mV (so 2 mHz coupling to the phase tracker). Under previous SNR conditions, our lines are typically at a few 100 Hz of amplitude, with a noise floor comparable to a few 100 mHz (SNR ~ 100s), so AM coupling seems not to be statistically limiting for 0.1% calibration.
Next:
- To further reject this residual coupling, it may be worth balancing the demodulated IQ amplitudes, by using the digital gains "C1:ALS-BEATX_FINE_Q_GAIN, C1:ALS-BEATX_FINE_I_GAIN".
- For now, this effect (and its frequency dependence above 300 Hz) can probably be neglected for ALS calibration purposes.
|
17396
|
Thu Jan 12 15:31:27 2023 |
Radhika | Update | ALS | XARM green laser lock debugging | [Radhika, Anchal, Paco]
AUX PDH Loop Stability
Today I tried aligning the XEND green beam into the arm cavity. Using M1 and M2 steering mirrors, I reached a max transmission ~1.2 of TEM00. In this configuration there was a "donut" mode also flashing, with transmission exceeding that of TEM00. Scanning all 4 degrees of freedom, I couldn't get TEM00 transmission to exceed 1.2, or significantly suppress the other modes. Not great mode matching. (PD gain: 20 dB; servo gain: 10.0.)
In an earlier conversation Paco had recommended I preamplify the green REFL signal with an SR560 before feeding it to the RF mixer. (For yarm this is done with an SR560 gain of 1000.) I did so and raised the gain on the SR560 until it overloaded (PD: 0 dB; SR560: 100). This didn't immediately improve the lock quality, but because alignment still needed work I wasn't surprised.
Anchal suggested the laser mode might be distorted by some lenses further upstream. We noticed some vertical spreading/distortion of the green beam by the first lens after SHG. I adjusted the pitch of an IR steering mirror until it disappeared. We then used the irises by the entrance to the arm cavity to coarsely align the input beam with M1 and M2. This time, fine alignment brought green transmission to just under 4. After slightly adjusting the half-wave plate, green transmission peaked at 4. (This is the highest I've seen it - previous max was 3.) The final combination of PD gain, SR560 gain, and servo gain that maximized transmission and duration of lock was (PD: 10 dB; SR560: 20; servo: 4.0). At its longest, lock on TEM00 was maintained for ~10 seconds.
AUX PDH Loop OLTF
In parallel with above, I was trying to take an OLTF of the loop whenever it was temporarily locked. I set up the measurement configuration like in the previous ELOG (injection at error point). Like last time, the loop would not lock when summing the PDH error signal with the excitation. I confirmed this was true even when I turned off the Moku excitation output. Checking the summed signal output, the Moku was adding an offset to the error signal. Buffering the excitation with an SR560 solved this issue.
The locked mode was switching pretty rapidly during the time I tried to measure the OLTF, and I ended up moving onto trying to improve lock. I might return today to try to take a measurement - I'll post it here. |
17395
|
Thu Jan 12 12:00:09 2023 |
Paco | Summary | LSC | FPMI BHD sensitivity curve with higher resolution upto 6.9 kHz | Here's the same sensitivity plot from yesterday (40m/17392), but with higher frequency resolution, upto 6.9 kHz, using GPS times from yesterday.
Now curves from FPMI with AS55_Q are also from yesterday, just before switching to BHD, so it will be more direct comparison |
17394
|
Thu Jan 12 10:06:10 2023 |
Paco | Update | ALS | DFD discriminant calibration | [Paco, Anchal] Log from yesterday work around 1Y2 rack; note that while this work was ongoing, TT2 position drifted slowly and misaligned the IFO input over the course of less than an hour. I suspect the DB9 breakout board and temporarily present components noted below may have introduced a floating ground in 1Y2, making the TT coil drivers misbehave. To support this claim, we noted that after removing the breakout board the drift disappeared!
We calibrated the DFD discriminant as a function of RF input level. The configuration was as follows:
- Break out IQ demod board RF output (in the rear chassis on 1Y2), looking at Ch1 board outputs (for BEATX). The two differential outputs were broken into two BNCs (pairs 1,6 and 2,7 for Q and I respectively), and fed into the Moku:Lab. A Marconi 2023A was used as a VCO, with FM ext mode enabled and a FreqDVN (modulation slope) of 200 kHZ / Vrms (1 Vrms = 1.41 Vp = 0.705 Vpp). The FM ext input on the Marconi was sourced by the OUT1 of Moku:Lab, and the IQ demod outputs were connected to IN1 and IN2 on Moku:Lab.
- After measuring the TF using a swept sine, I verified that the frequency response was flat up to 150 kHz (Marconi cuts FM ext at 275 kHz), so I switched to the oscilloscope instrument and setup OUTPUT to a 211.1 Hz sine wave, 1.41 Vpp to dither the 40 MHz by +- 100 kHz.
- Using the arbitrary math function in Moku:Lab Scope, I computed the DFD output magnitude = sqrt(I**2 + Q**2), and measured its mean over 3 seconds.
The results are summarized in Attachment #1. |
17393
|
Wed Jan 11 17:05:55 2023 |
Anchal | Update | ASC | WFS demodulation board 112B - Working as expected | The other modified board 112B has been fixed and tested now. See the results attached. The issue was in some malfunctioning OP284 which have been replaced by AD8672. |
17392
|
Wed Jan 11 16:56:57 2023 |
yuta | Summary | LSC | FPMI BHD recovered, LO phase noise not limiting the sensitivity | [Paco, Yuta]
We recovered FPMI BHD, and sensitivity was estimated. High frequency sensitivity is improved by an order of magnitude compared with AS55 FPMI.
We also estimated the contribution from LO phase noise, and found that LO phase noise is not limiting the sensitivity.
Locking sequence:
1. Lock electronic FPMI
DARM:
- 0.5 * POX11_I - 0.5 * POY11_I
- DARM filter module, FM4,5 for acquisition, FM1,2,3,6,8 triggered, C1:LSC-DARM_GAIN = 0.015 (gain lowered from BHD FPMI in December (was 0.02) to have more gain margin)
- Actuation on 0.5 * ETMX - 0.5 * ETMY
- UGF ~ 150 Hz
CARM:
- 0.5 * POX11_I + 0.5 * POY11_I
- CARM filter module, FM4,5 for acquisition, FM1,2,3,6,8 triggered, C1:LSC-CARM_GAIN = 0.012
- Actuation on -0.734 * MC2
2. Lock MICH
MICH:
- 1.1 * REFL55_Q
- MICH filter module, FM4,5,8 for acquisition, FM2,3,6 triggered, C1:LSC-MICH_GAIN = +10
- Actuation on 0.5 * BS
- UGF ~35 Hz (Attachment #1)
3. Hand over to real DARM/CARM
DARM:
- 2.617 * AS55_Q
CARM:
- 0.496 * REFL55_I
- UGF ~200 Hz (Attachment #2)
4. Lock LO_PHASE
LO_PHASE:
- 1 * BH55_Q (demod phase: -110 deg; this was chosen by hand to maximize fringe in BH55_Q when FPMI is locked. This seemed to make more robust LO_PHASE lock, compared with December)
- FM5,8, C1:HPC-LO_PHASE_GAIN=-1
- Acuation on 1 * LO1
- UGF ~110 Hz (Attachment #3)
5. Hand over to BHD_DIFF
DARM:
- -1.91 * BHD_DIFF (ratio between AS55_Q to BHD_DIFF was measured with a DARM line at 575.125 Hz, which was measured to be -0.455; BHD_DIFF was zeroed by balancing A and B before the measurement)
- UGF ~150Hz (Attachment #4)
Measured sensing matrix:
/opt/rtcds/caltech/c1/Git/40m/scripts/CAL/SensingMatrix/ReadSensMat.ipynb
Sensing matrix with the following demodulation phases (counts/counts)
{'AS55': -168.5, 'REFL55': 92.32, 'BH55': -110.0}
Sensors DARM @307.88 Hz CARM @309.21 Hz MICH @311.1 Hz LO1 @315.17 Hz
AS55_I (-0.35+/-1.44)e-02 (+3.23+/-1.09)e-02 (-0.18+/-9.84)e-03 (+0.02+/-1.33)e-03
AS55_Q (-0.45+/-3.48)e-02 (-0.19+/-1.23)e-02 (+0.05+/-1.43)e-03 (-0.02+/-2.03)e-04
REFL55_I (+0.05+/-1.31)e-01 (+4.51+/-0.07)e-01 (-0.03+/-3.44)e-02 (+0.11+/-1.01)e-03
REFL55_Q (-0.39+/-4.74)e-04 (-1.36+/-0.50)e-03 (+0.16+/-2.80)e-04 (+0.10+/-4.92)e-05
BH55_I (-1.90+/-5.00)e-03 (+0.80+/-8.82)e-03 (-0.50+/-2.29)e-03 (-4.61+/-9.52)e-04
BH55_Q (-0.31+/-4.86)e-02 (-3.13+/-3.04)e-02 (-0.06+/-1.07)e-02 (-1.42+/-2.11)e-03
BHDC_DIFF (-5.56+/-0.21)e-02 (+0.10+/-1.64)e-02 (+0.10+/-1.40)e-03 (+0.77+/-2.27)e-04
BHDC_SUM (+1.75+/-3.90)e-03 (-1.22+/-3.22)e-03 (-0.35+/-1.52)e-03 (-0.36+/-6.42)e-04
Sensing matrix with the following demodulation phases (counts/m)
{'AS55': -168.5, 'REFL55': 92.32, 'BH55': -110.0}
Sensors DARM @307.88 Hz CARM @309.21 Hz MICH @311.1 Hz LO1 @315.17 Hz
AS55_I (-0.30+/-1.25)e+11 (+2.18+/-0.73)e+11 (-0.07+/-3.59)e+10 (+0.08+/-5.02)e+09
AS55_Q (-0.39+/-3.03)e+11 (-1.27+/-8.32)e+10 (+0.20+/-5.23)e+09 (-0.06+/-7.66)e+08
REFL55_I (+0.04+/-1.14)e+12 (+3.05+/-0.05)e+12 (-0.01+/-1.26)e+11 (+0.42+/-3.82)e+09
REFL55_Q (-0.34+/-4.12)e+09 (-9.20+/-3.37)e+09 (+0.06+/-1.02)e+09 (+0.04+/-1.86)e+08
BH55_I (-1.65+/-4.34)e+10 (+0.54+/-5.95)e+10 (-1.83+/-8.36)e+09 (-1.74+/-3.59)e+09
BH55_Q (-0.27+/-4.23)e+11 (-2.11+/-2.05)e+11 (-0.21+/-3.91)e+10 (-5.36+/-7.98)e+09
BHDC_DIFF (-4.83+/-0.18)e+11 (+0.07+/-1.11)e+11 (+0.35+/-5.09)e+09 (+2.90+/-8.56)e+08
BHDC_SUM (+1.52+/-3.39)e+10 (-0.82+/-2.17)e+10 (-1.29+/-5.56)e+09 (-0.14+/-2.42)e+09
NOTE that some of sensing matrix element (e.g. DARM to AS55_Q) is wrong, because of ill defined sign in C1:CAL-SENSMAT_XXXX_XXXX_AMPMON channels.
Locked GPS times:
- 1357511737 to 1357513050
- 1357513448 to 1357518188 (intentionally unlocked)
Sensitivity estimate:
- See Attachment #5 and #6 (high frequency zoomed), dashed traces with darker colors are of AS55_Q FPMI from 40m/17369.
- DARM_IN1 was calibrated using DARM content in BHD_DIFF, with 1 / (1.91 * 4.83e11 counts/m) = 1.086e-12 m/counts (which should be similar to DARM_IN1 calibration for AS55_Q, because we are balancing the error signals going to DARM_IN1, and it is as expected; see 40m/17369).
- DARM_OUT calibration is the same as 40m/17369; ETM plant = 10.91e-9 / f^2 m/counts.
- LO phase noise was estimated using BH55_Q, with the collowing calibration factor (BH55_Q calibrated into LO1 motion, into BHD_DIFF, and then into DARM).
1/5.36e9*2.9e8/4.83e11 = 1.15e-13 m/counts
- Seismic noise was estimated using C1:IOO-MC_F_DQ, with the same calibration factor found in 40m/16975.
- Dark noise was estimated using DARM_IN1 when the PSL shutter is closed.
Discussion:
- Sensitivity below ~10 Hz is probably limited by seismic noise
- Noise above 1 kHz might be limited by dark noise of BHD_DIFF, but not below 1 kHz. For AS55_Q FPMI, sensitivity above ~300 Hz was limited by dark noise.
- LO phase noise is not limiting the sensitivity, from the estimated noise using BH55_Q.
- Both BH55_Q and BHD_DIFF have funny structure like forest above ~100 Hz. This might be from suspensions in the AS path and LO path to BHD. It could also be that calibration lines for measuring sensing matrix were too much (BHD FPMI sensitivity was measured with calibration lines around ~310 Hz on).
Next:
- Update the c1cal model to put correct signs to C1:CAL-SENSMAT_XXXX_XXXX_AMPMON channels.
- Measure BHD FPMI sensitivity with calibration lines off.
- Find better LO phase to improve the sensitivity.
- Lock LO phase with RF+audio and RF44 and compare the sensitivity.
- Move on to PRFPMI BHD. |
17391
|
Tue Jan 10 20:24:29 2023 |
Anchal | Update | ASC | WFS demodulation board 111B - Working as expected | I've completed the modifications on two WFS demod boards. This required replacing all 8 mixer ICs on each board. I also tuned each channel to get less than 2 mV offset on all of them.
I was able to complete testing the board SNo. 111B today. The results are attached. The test was done by feeding the board 22 MHz LO generated by frequency doubling. A signal at 11 MHz was generated using Moku:Lab at 1mVpp and then further attenuated by 10 dB to make a fair comparison with the previous testing of the IMC WFS board at 29.5 MHz. This board has the same response as the IMC WFS board at 29.5 MHz. I tested all four channels in the second plot.
I'll complete the testing of the second board SNo 112 B and then move on to setting up the optical path for AS WFS. |
17390
|
Tue Jan 10 16:06:58 2023 |
yuta | Summary | PSL | PMC transmission dropped to 0.68 | [JC, Paco, Yuta]
It seems like PMC transmission (C1:PSL-PMC_PMCTRANSPD) dropped to ~0.68 from ~0.74 on Dec 27.
We tried to tweak PZT offset for PMC loop and input alignment to PMC, but PMC transmission didn't increased.
PSL laser temperature was also sweeped in the range 30.3 - 31.6 degC, but didn't help. The PSL temperature was reverted to original 30.61(1) degC.
Power measured at PSL output now is 893 mW (measured at our standard place shown in 40m/16672), which used to be 951 mW in June 2022 (40m/16886).
Power measured at PMC input (see attached photo) now is 1.18 W.
Next:
- What was the previous PMC input power we had?
- Sweep PSL laser temperature for larger range |
17389
|
Tue Jan 10 15:07:33 2023 |
JC | Update | General | Clean Large Optical Table | I have been working on cleaning the large optical table next to the PSL table. I have placed the known optics along in the optical cabinets Section Y5. Post and pillars have been placed in a cabinet along X-Arm. (I will work on gathering optics and optical hardware into a single area in the future, having them separated is only temporary.) I found some optics that were chipped and/or unlabeled. These have been placed in a box.
There were cables coming into this table from the wire rack above, I plan on removing this. If there are any issues with me doing this, please message me.
My main purpose for cleaning this is to potentially have an open area to place OMC for the next vent. We can also move the PD Testing table here. |
17388
|
Mon Jan 9 19:41:01 2023 |
Anchal | Update | SUS | Null stream (butterfly/pringle) row added and DQed | I updated the suspension model (/cvs/cds/rtcds/userapps/trunk/sus/c1/models/lib/sus_single_control_new.dml) to add a 5th row in the input matrix so that we can put in the calculated NULL stream vector (also have been called as Butterfly mode or Pringle mode in the past). The output of this row would go through a filter module and then is sent to the testpoint named 'C1:SUS-OPT_SENSOR_NULL' where OPT is the optic acronym. This channel is acquired in frames at 256 Hz and would be available as C1:SUS-OPT_SENSOR_NULL_DQ. After the update in the model, I built, installed and restarted models c1sus, c1mcs, c1scx, c1scy, c1su2, and c1su3. Then I restarted daqd, and everything came up nicely. After but restore, I added the null stream vector for the optics it was already known from a free swing test. ITMX, ITMY, ETMX, PRM, and SRM null stream vectors needs to be calculated from the other 4 rows. It is set to zero right now. Medm screen for the input matrix was also updated to allow seeing this row.
|
17387
|
Thu Jan 5 14:30:32 2023 |
Paco | HowTo | DAQ | nds2 server restart | After being unable to fetch data offsite using the nds40 server, I found enlightment here. Our nds2 server is running on megatron and the link before should be sufficient to restore it after any hiccups.  |
17386
|
Wed Jan 4 17:15:41 2023 |
Koji | Configuration | CDS | unknown dhcp request to fb1 | The dhcpd error on the log file stopped when the yellow (DAQ) ethernet cable was removed. With Chris's permission I left it unconnected (Attachment 1).
Chris pinted that the IPMI on c1shimmer is supposed to be exposed to 192.168.113. net rather than DAQ net.
From dhcpd.conf on chiara:
host c1shimmer-ipmi {
hardware ethernet 3c:ec:ef:c8:44:78;
fixed-address 192.168.113.37;
}
So the ethernet connections of c1shimmer is still questionable. The next person to work with c1shimmer needs to check them.
This would also be related?
https://lanforge.wordpress.com/2015/11/10/turning-off-ipmi-dhcp/ |
17385
|
Wed Jan 4 15:10:19 2023 |
Koji | Configuration | CDS | unknown dhcp request to fb1 | Jamie reported that:
The logs (/var/log/daemon.log) on fb1 are filling with this line:
Jan 03 14:11:51 fb1 dhcpd[1152]: DHCPDISCOVER from 3c:ec:ef:c8:44:78 via enp3s0: network 10.0.113.0/24: no free leases
It seems that some machine on the network is trying to get an IP address but can't
- The MAC address 3c:ec:ef:xx:xx:xx indicates this is one of the supermicro units.
- The IP address indicates this is on the DAQ network which fb1 is spanning.
- There is a switch for this DAQ network.
- There are 8 machines connected to the switch. fb1 is via optical, and the other 7 have yellow ethernet cables (Attachment 1).
- fb1 and other 6 RT machines already have 10.0.113.x assigned.
- The rest is c1shimmer. I can't ssh into it from martian. KVM on rack 1X7 shows a black screen assuming 1-7 is c1shimmer. I wonder what's the status of this machine, but left untouched so far.
|
17384
|
Wed Jan 4 12:12:57 2023 |
Paco | Summary | Calibration | ALS calibration error from DFD | [Paco, Anchal]
One of the crucial and currently limiting calibration errors is in the ALS beat. We think a major driver of calibration uncertainty may be the DFD TF calibration since it depends on the RF beat power.
The RF beat power as monitored by C1:ALS-BEATY_RF_POW shows relative fluctuations of 0.02% at the 16 Hz sampling rate (actually 0.09% rms from Attachment #1) once the single arm and YAUX laser are both locked. This sounds ok, but that's just a statistical estimate. The beat frequency changes every time we break and reacquire the lock, making the BEAT PD frequency response and DFD overall calibration change their nominal TF. This changes can be significantly larger than 0.02% (e.g 4.9% = observed change in the RF power after breaking and reacquiring the YAUX lock this morning). This is far more significant than the contribution from the RIN on the two lasers.
This kind of systematic error could be addressed by either/all:
- Stabilizing the ALS BEAT absolute frequency (offset locking using freq counter and simple integrator or equivalent)
- Rejecting RF beat amplitude fluctuations by mixing a stable DC voltage in the DFD, or using a comparator.
- Our plan this afternoon is to quickly measure the RF AM rejection level from a simple mixer and a stable voltage reference and plan ahead.
- ISS on both lasers
|
17383
|
Wed Jan 4 08:52:11 2023 |
JC | Update | OPLEV Tables | ETMX laser has died | Koji mentioned to me that this laser has a degraded output. I removed this replacement laser and put in a new laser from the Y Arm Cabinets, Attachment #1. I placed the degraded laser back to its visitor setup and tested with the power supply, this is shown in Attachment #2. Along with this, I also labeled this laser as degraded. |
17382
|
Tue Jan 3 17:41:55 2023 |
Koji | Update | CDS | Chiara local backup | Chris modified the fstab so that the disk is automatically mounted. The key was that these two disks have an identical UUID and use this UUIS for mounting. I confirmed that /cvs/cds was properly backup-ed on Jan 4th.
controls@chiara:~$ lsblk
NAME MAJ:MIN RM SIZE RO TYPE MOUNTPOINT
sda 8:0 0 1.8T 0 disk
├─sda1 8:1 0 512M 0 part /boot/efi
├─sda2 8:2 0 1.8T 0 part /
└─sda3 8:3 0 31.9G 0 part [SWAP]
sdb 8:16 0 5.5T 0 disk
└─sdb1 8:17 0 5.5T 0 part /home/cds
sdc 8:32 0 5.5T 0 disk
└─sdc1 8:33 0 5.5T 0 part
└─md0 9:0 0 5.5T 0 raid1 /media/40mBackup
sdd 8:48 0 5.5T 0 disk
└─sdd1 8:49 0 5.5T 0 part
└─md0 9:0 0 5.5T 0 raid1 /media/40mBackup
controls@chiara:~$ sudo blkid
/dev/sdc1: UUID="052f9129-f2eb-6ef1-473d-a748a1ffa5d3" UUID_SUB="9ac96976-e1d5-cc01-13f7-10944d5d6735" LABEL="chiara:0" TYPE="linux_raid_member" PARTLABEL="primary" PARTUUID="5cf6aac7-3f29-42e0-8db6-de2978ca84f0"
/dev/sdd1: UUID="052f9129-f2eb-6ef1-473d-a748a1ffa5d3" UUID_SUB="9f124649-ff79-e406-54ab-6374b4acddf1" LABEL="chiara:0" TYPE="linux_raid_member" PARTLABEL="primary" PARTUUID="01296513-46cf-4210-8d59-6f6bc25934ee"
/dev/sda1: UUID="69C0-DD2F" TYPE="vfat" PARTUUID="00ac6b04-2675-4842-9193-ee5c4575b4b4"
/dev/sda2: UUID="75f19654-1504-4548-b9a7-95f22d507cb4" TYPE="ext4" PARTUUID="f5022aff-7f84-4e72-9dfa-44b0412e907b"
/dev/sda3: UUID="e9d6a5b7-78e7-4373-9bac-0509458860cd" TYPE="swap" PARTUUID="f7388c20-b17e-4f21-9093-1df58d00d9e3"
/dev/sdb1: LABEL="cvs" UUID="eceb000f-9493-45fd-85f1-2c95c6819f4a" TYPE="ext4" PARTUUID="2e090790-7d03-406d-8ee9-5a5c96e14258"
/dev/md0: UUID="6d16f518-1332-42c3-ad16-0a9d384c6dad" TYPE="ext4"
controls@chiara:~$ cat /etc/fstab
# /etc/fstab: static file system information.
#
# Use 'blkid' to print the universally unique identifier for a
# device; this may be used with UUID= as a more robust way to name devices
# that works even if disks are added and removed. See fstab(5).
#
# <file system> <mount point> <type> <options> <dump> <pass>
# / was on /dev/sda2 during installation
UUID=75f19654-1504-4548-b9a7-95f22d507cb4 / ext4 errors=remount-ro 0 1
# /boot/efi was on /dev/sda1 during installation
UUID=69C0-DD2F /boot/efi vfat defaults 0 0
# swap was on /dev/sda3 during installation
UUID=e9d6a5b7-78e7-4373-9bac-0509458860cd none swap sw 0 0
UUID=eceb000f-9493-45fd-85f1-2c95c6819f4a /home/cds ext4 errors=remount-ro 0 1
UUID=6d16f518-1332-42c3-ad16-0a9d384c6dad /media/40mBackup ext4 errors=remount-ro 0 1
|
17381
|
Tue Jan 3 16:11:05 2023 |
Koji | Update | CDS | Chiara local backup | I checked how the local backup of our /cvs/cds is done. The last backup was taken on 2022-12-15 and it kept failing since 2022-12-16.
controls@chiara:/opt/rtcds/caltech/c1/scripts/backup$ tail -100 localbackup.log
....
2022-12-15 07:00:02,404 INFO Updating backup image of /cvs/cds
2022-12-15 07:17:51,901 INFO Backup rsync job ran successfully, transferred 1,352 (reg: 1,288, dir: 64) files.
2022-12-16 07:00:01,650 INFO Updating backup image of /cvs/cds
2022-12-16 07:00:01,668 ERROR External drive not mounted!!!
....
The backup destination is /media/40mBackup according to /opt/rtcds/caltech/c1/scripts/backup/localbackup
During the power outage on Dec 15, 2022 (<- not well elogged), the volume was probably unmounted.
This disk is missing although some spare disks exist. The disks are indicated as RAID1. I'll ask Chris if he knows what the current configuration is and how to add it to fstab.
controls@chiara:/opt/rtcds/caltech/c1/scripts/backup$ df
Filesystem 1K-blocks Used Available Use% Mounted on
udev 16394928 0 16394928 0% /dev
tmpfs 3283044 123944 3159100 4% /run
/dev/sda2 1888292368 13161244 1779137424 1% /
tmpfs 16415220 0 16415220 0% /dev/shm
tmpfs 5120 0 5120 0% /run/lock
tmpfs 16415220 0 16415220 0% /sys/fs/cgroup
/dev/sda1 523248 3620 519628 1% /boot/efi
/dev/sdb1 5814157460 2604230900 2916884052 48% /home/cds
tmpfs 3283044 8 3283036 1% /run/user/1001
tmpfs 3283044 4 3283040 1% /run/user/114
controls@chiara:/opt/rtcds/caltech/c1/scripts/backup$ lsblk
NAME MAJ:MIN RM SIZE RO TYPE MOUNTPOINT
sda 8:0 0 1.8T 0 disk
├─sda1 8:1 0 512M 0 part /boot/efi
├─sda2 8:2 0 1.8T 0 part /
└─sda3 8:3 0 31.9G 0 part [SWAP]
sdb 8:16 0 5.5T 0 disk
└─sdb1 8:17 0 5.5T 0 part /home/cds
sdc 8:32 0 5.5T 0 disk
└─sdc1 8:33 0 5.5T 0 part
└─md0 9:0 0 5.5T 0 raid1
sdd 8:48 0 5.5T 0 disk
└─sdd1 8:49 0 5.5T 0 part
└─md0 9:0 0 5.5T 0 raid1
|
17380
|
Tue Jan 3 15:28:12 2023 |
JC | Update | OPLEV Tables | ETMX laser has died | Seems that the ETMX laser died over the break and took an extended vacation. I swapped the laser out with one Koji would use to model Suspension wires.
The laser which died out seems a bit new though, it was manufactured in Jan 2019. A photo of this laser is attached. |
|